an object thrown up from a cliff at 10m/s reaches velocity of 20m/s down as it lands , if the acceleration due to gravity is 9,8m/s, what is the object displacement


1
Expert's answer
2021-05-12T14:58:50-0400

Let's take the upwards as the positive direction. Then, we can find the object displacement from the kinematic equation:


"v^2=v_0^2+2gy,""y=\dfrac{v^2-v_0^2}{2g},""y=\dfrac{(-20\ \dfrac{m}{s})^2-(10\ \dfrac{m}{s})^2}{2\cdot(-9.8\ \dfrac{m}{s^2})}=-15.31\ m."